OpenStudy (anonymous):

suppose 100 radioactive atoms have a half life of 10 seconds. Approximately how many radioactive atoms will be left after 20 seconds. 50 12.5 25 75

OpenStudy (anonymous):

25 as two half lives are happening :D 1/2 * 1/2 = 1/4 1/4 * 100 = 25

OpenStudy (anonymous):

half life means the time taken for the amount of a substance undergoing decay to decrease by half its initial value, so in 10 seconds it decreases from 100 to 50, in 20 it decreases from 50 to 25